From The Munich Maker Lab's Wiki
Jump to navigation Jump to search


Date User State Notes
Jul 18 Robert and various πŸ›‘ Disassembly and cleanup. Preparations for new laser Tube


Date User State Notes
Jul 10 PeterF and various βœ… Laser cutter was setup, calibrated, the exhaust system was modified and the Laser PC was reinstalled with Win10


Date User State Notes
Dec 21 AdrianM πŸ›‘ We're moving currently, so the Laser is packed up in the corner.
Feb 10 Bernhard βœ… Bernhard realigned the tube (again) and it should work as desired
Feb 3rd Severin πŸ›‘ Laser out of order, not to be used. No real power observed by Bernhard when trying to cut things. The last object (woodern spatula) engraved when the laser broke is hanging on the pegboard next to the vending machine.


Date User State Notes
Dec 29th David βœ… Beam Combiner installed and calibration of tube, mirrors and combiner done. Working fine on the whole cutting area. Installed lens: 75mm
Dec 27th David πŸ›‘ DO NOT USE! Complete recalibration necessary! In the course of this recalibration a complete rework of the rear laser-combiner-mirror assembly is planned.
May 11th David βœ… Laser was calibrated and filters were replaced. Working for now, but a complete overhaul is necessary. Installed lens: 75mm
Mar 28th Severin πŸ›‘ Can't cut through 5mm MDF; 3mm MDF needs way to much power -> probably needs recalibration. Do not use until then.
Jan 28th Eax βœ… Installed the 75mm lens. Put a sign indicating the installed lens on the top cover of the Cutter.
Jan 14th Eax βœ… Laser IN order! Housing was closed, replacement part were cut for future maintenance.
Jan 9th Eax πŸ›‘ Laser out of order, not to be used. Laser mirrors are adjusted and calibrated. Housing needs to be closed. Beam combiner needs to be mounted and positioned correctly.


Date User State Notes
Dec 23rd Eax πŸ›‘ Laser out of order, not to be used. The drift has been corrected by balancing the position of the mirrors and the linear rails. What still remains to be done is the *angular* alignment of the mirrors and the laser tube. The positional alignment of the mirrors has been corrected at this time. The first mirror was assembled 8mm to high for the first linear rail, and there was a misalignment of 3mm in the horizontal direction.
Dec 1st Eax πŸ›‘ Laser out of order, not to be used. Mirror cover on the laser tube itself was loose, because presumably s.o. tried to attach a too tight pilot laser holding assembly. Fixed the mirror cover to its place by using glue.
Nov 29 Eax πŸ›‘ Laser out of order, not to be used. Replaced all three mirrors with the new batch of mirrors. Determined that the position of the 1st mirror is not aligned with the linear axis of the 2nd mirror (X axis movement). This leads to a drift of the spot on the 2nd mirror. The position of the first mirror needs correction. For this a pilot laser (red visible) would be ideal.
Nov 2nd Tiefpunkt πŸ›‘ Laser out of order, not to be used. Mirror 2 was loose again, after retightning not really alligned. Before realigning, mirrors should be switched for the now ones, which should fit the mirror mounts much better. Peter tried some calibration, but one of the screws of mirror 2 was too tight and the plastic handle on it broke. Reattached with superglue.
Oct 19th Siedi βœ… Replaced SSR for the exhaust air. The switch is working again.
Oct 18th Eax βœ… Laser cutter is operable. Air compressor was removed from the frame. Water protect issue not reproducible. 75mm lens assembled. 50mm lens is missing.
Oct 14th Eax βœ… Laser is calibrated. BUT: Air compressor not working AND "Water protect" triggered every now and then.
Oct 7th Tiefpunkt (talk) ⚠️ SSR of exhaust broken, won't turn exhaust off. Severin will get a replacement part.
Oct 5th Tiefpunkt (talk) πŸ›‘ Laser out of alignment again after longer laser job. Mirror assembly no. 2 gets warm, and loosens screw again. Repeated hot-cold cycles will make it the screw loose and the mirror wiggle out of position. Laser deemed not usable until realignment incl. mirror no. 1 as well. We'll look for a replacement mirror as well.
Sep 30th Tiefpunkt (talk) βœ… Laser out of alignment again. Mirror No. 2 loose inside its holder. Mirror assembly screwed out of holder, mirror cleaned, and put back into assembly. Assembly holds 25mm mirrors, but the mirror seems slightly bigger, thus does not sit flush in the assembly, just jammed in there as tight as possible. Mirror put back in, and mirror no. 2 realigned, works fairly well for now.
Sep 28th Tiefpunkt (talk) βœ… Laser aligned by Peter